### v1

# DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

Abstract
Native multimodal large language models (MLLMs) restructure a single large language model (LLM) into a spoken language model (SLM) capable of both speech and text generation. Compared to modular and aligned MLLMs, native MLLMs preserve richer paralinguistic features such as emotion and prosody, and generate speech responses directly within the backbone LLM rather than using a separate speech decoder. This integration also results in lower response latency and smoother interaction. However, native MLLMs suffer from catastrophic forgetting and performance degradation because the available paired speech-text data is insufficient to support the pretraining of MLLMs compared to the vast amount of text data required to pretrain text LLMs. To address this issue, we propose DeepTalk, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs. Meanwhile, the end-to-end dialogue latency remains within $0.5$ seconds, ensuring a seamless and intelligent speech interaction experience. Code and models are released at https://github.com/talkking/DeepTalk.
1 Introduction
Since GPT-4o GPT4o has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models. As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) radford2023robust ; bai2024seed ; povey2011kaldi , the resulting text is then comprehended by a large language model (LLM) liu2024deepseek ; yang2024qwen2 ; glm2024chatglm to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) du2024cosyvoice ; ren2019fastspeech ; wang2025spark model. However, this cascaded structure results in high latency, and errors from each module can accumulate and propagate, leading to degraded performance. Therefore, recent speech LLMs adopt an end-to-end approach, using a unified model to process both speech and text simultaneously, which enhances the capabilities of speech understanding and generation. At the same time, this approach significantly reduces interaction latency and provides a more natural and smooth voice interaction experience.
Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal chen2025minmo . Representatives of the former include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , Freeze-Omni wang2024freeze , and VITA-1.5 fu2025vita . These models connect a speech encoder and a speech decoder to the LLM to handle audio understanding and audio generation tasks separately, placing semantic understanding within the LLM module. This maximally the retains LLM’s general capabilities. Furthermore, since the LLM itself does not require retraining, much less speech-text paired pretraining data is needed.
Representative of the latter including Mini-Omni xie2024mini , GLM-4-Voice zeng2024glm , Moshi defossez2024moshi , and LUCY gao2025lucy use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens. Compared to modular aligned multimodal approaches, native ones greatly reduces the latency of speech token output. More paralinguistic information in speech, such as emotion and prosody, is retained within the SLM rather than in the speech decoder, benefiting the full expression of paralinguistic information. However, since native multimodal models expand the original LLM’s vocabulary and require retraining, they need a large amount of audio-text paired pretraining data. Therefore, in real-world scenarios where audio data is much less abundant than text data, native multimodal models are prone to catastrophic forgetting, severely impairing the original LLM’s language capabilities.
In this paper, to alleviate the catastrophic forgetting faced by native multimodal models zhai2023investigating , we propose DeepTalk, a model that applies the mixture of experts (MoE) for multimodal learning li2025uni ; li2024cumo ; zhong2024moextend ; shen2024mome to effectively isolate modality-specific knowledge. Since most speech dialogue data are conversational and colloquial, while text LLMs are primarily trained on formal written text, training LLMs with colloquial speech dialogue data can lead to less standardized written outputs, resulting in decreased language performance on certain LLM benchmarks. Therefore, we assign different parameters within the LLM to specialize in instruction data from different modalities, allowing modality experts to focus on single-modality training and avoid interference between modalities. Finally, we use cross-modal instruction data to jointly train the modality experts, enabling the model to output both speech and text simultaneously.
To minimize the damage to the original text LLM, we propose an adaptive modality expert selection strategy, which dynamically selects modality experts based on the MoE model’s modality load on different data. Experts with a high speech token load but a low text token load are selected as audio experts. After several iterations of modality expert selection and training, DeepTalk achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
In summary, the contributions of this paper are as follows:
  • We propose DeepTalk, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge. At the same time, modalities can collaborate to jointly output multimodal results, enabling both colloquial and formal language to coexist within a single model.
  • We introduce an adaptive modality expert selection method based on modality load, which further reduces the damage to the original text LLM and alleviates the problem of catastrophic forgetting.
  • To the best of our knowledge, this is the first native multimodal large speech interaction model based on the MoE architecture. It opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
Figure 1: Comparison of DeepTalk and existing voice interaction systems. Compared to modular MLLMs, DeepTalk embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM. Through adaptive modality expert selection and a three-stage training process for modality experts, DeepTalk effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
2 Related Work
2.1 End-to-End Speech Interaction System
Speech is among the most important signal in human communication, as it not only conveys semantic content but also carries various paralinguistic information such as emotion, speech rate, intonation, and timbre. As a result, speech interaction has become an increasingly important mode of human-computer interaction. Traditional speech interaction systems are typically composed of a pipeline of separately optimized ASR, LLM, and TTS modules. Such systems not only suffer from high interaction latency, but also accumulate errors from each module, leading to a less satisfactory overall user experience. Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach ji2024wavchat .
Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM kong2020panns ; chu2024qwen2 ; chu2023qwen ; das2024speechverse . This approach not only provides the LLM with richer speech information than using pure text input, but also reduces the latency caused by ASR recognition. However, these models still require an additional TTS module to achieve end-to-end speech interaction. To provide rich speech information, LLaMA-Omni fang2024llama feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens. Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss graves2006connectionist is used to align the upsampled hidden states with the audio token outputs. Meanwhile, to reduce latency for real-time speech interaction, a predefined chunk size is set to enable the vocoder to synthesize speech in a streaming manner. Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of $82$Hz and adopts a delayed decoding strategy copet2023simple ; peng2024voicecraft ; lyth2024natural to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference. Minmo chen2025minmo integrates a language model with an autoregressive CosyVoice2 decoder du2024cosyvoice . The hidden states of the language model are fed into the CosyVoice2 decoder to autoregressively predict speech tokens, which are then passed to a vocoder for waveform synthesis.
Fully end-to-end approaches integrate the TTS module into the LLM, allowing the LLM itself to generate both text and speech tokens, rather than relying on a separate speech decoder or TTS module for speech synthesis. Discrete speech tokens are typically generated by neural audio codecs defossez2022high ; ji2024wavtokenizer ; zhang2023speechtokenizer or self-supervised models lakhotia2021generative ; hsu2021hubert ; chen2022wavlm . Currently, there are two mainstream modeling paradigm for end-to-end speech systems. The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice zeng2024glm and Spirit-LM nguyen2025spirit , where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens. The second is parallel audio-text modeling, represented by models such as Mini-Omni xie2024mini , Moshi defossez2024moshi , Slam-Omni chen2024slam and LUCY gao2025lucy . In this paradigm, text and audio tokens can be generated in parallel: while generating text tokens, multiple LM heads are used to simultaneously predict multiple audio tokens. Compared to the interleaved approach, parallel modeling compresses speech tokens into shorter sequences, enabling the modeling of higher-bitrate speech tokens.
Modular Speech Language Models  Modular large speech models are primarily built on an architecture where a LLM is directly connected to an audio encoder and decoder through adapters. Examples include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , and Freeze-Omni wang2024freeze , as shown in Fig. 1 b). In this structure, the LLM is only responsible for textual decoding, while speech decoding is handled by the audio decoder. As a result, the language capabilities of the LLM are largely preserved. Additionally, Minmo employs Low-Rank Adaptation (LoRA) hu2022lora for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities. However, such modular speech-language models face several challenges. For instance, the transmission of speech information relies entirely on the audio decoder, with the LLM only responsible for conveying textual content. This means that the expression of paralinguistic information in speech is completely dependent on the speech decoder, resulting in underutilization of the LLM. Furthermore, there are additional issues such as low deployment efficiency and the complexity of balancing the capacities of the LLM and the SLM.
Native Speech Language Models Native speech language models reconstruct the LLM into an SLM, enabling the LLM to output both text and speech tokens simultaneously, as shown in Fig. 1 c). Currently, most mainstream native multimodal approaches such as Mini-Omni xie2024mini , Moshi defossez2024moshi , GLM-4-Voice zeng2024glm , and LUCY gao2025lucy have the LLM handle both text and speech decoding. This approach offers low latency, easy deployment, and allows a unified model to process both audio and text decoding, which better aligns with end-to-end requirements. However, without large-scale speech-text paired data, retraining the LLM in this way can lead to catastrophic forgetting and performance degradation.
2.2 Multimodal Mixture-of-Experts
BEiT-3 wang2023image , VLMo bao2022vlmo , and Uni-MoE li2025uni employ specific modality expert groups to capture modality-specific information. MoExtend zhong2024moextend further expands modality experts horizontally on top of existing LLMs to capture additional modality information. CuMo li2024cumo extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME shen2024mome uses multiple feature encoders as modality expert groups to encode multi-dimensional features. VL-MoE shen2023scaling and MoE-LLaVa lin2024moe introduce mixture-of-experts (MoE) to improve training and deployment. MoMa lin2024moma pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens. Inspired by these works, we introduce multimodal mixture-of-experts (speech experts and text experts) into end-to-end large speech interaction models for native MLLM pretraining. At the same time, we use the modality load of each expert to adaptively select modality experts, a method called Adaptive Modality-Specific MoE, to address the catastrophic forgetting problem aforementioned.
3 Methods
3.1 Architecture
The overall architecture of the model is shown in Fig. 2, which mainly consists of an Audio Encoder, a connector, an MoE backbone, and a streaming codec. Let $X_{i}^{A}$ and $X_{i}^{T}$denote the audio input and text input of the $i$-th sample, respectively, and $Y_{i}^{A}$ and $Y_{i}^{T}$ denote the audio output and text output of the $i$-th sample, respectively. The model’s text and speech responses are represented as $Y^{T}\in V$ and $Y^{A}\in U$, where $V$ and $U$ denote the text vocabulary and the speech codec vocabulary, respectively. The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model, as shown below:
$H_{i}^{A}=Adapter(Encoder(X_{i}^{A})),$ (1)
$H_{i}^{T}=Embedding(X_{i}^{T}),$ (2)
$H_{i}=mean(H_{i}^{A},H_{i}^{T}),$ (3)
where $mean$ here refers to taking the average of the two embeddings. The loss L of the model over N samples can be defined as:
$\textit{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}logP(Y_{i,t}^{T},Y_{i,t}^{A}|Y_{i,<t}^{T},Y_{i,<t}^{A},H_{i}),$ (4)
where $T_{i}$ donates the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
Audio Encoder and Adapter We use Whisper-medium radford2023robust as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads. It includes multiple convolutional downsampling layers that downsample the speech features by a factor of 4. We use an 80-dimensional log Mel-filterbank with a 25ms window length computed every 10ms as the input to the audio encoder. A convolution-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features. Finally, for the input $X_{i}^{A}$, as shown in Equation 1, we obtain the hidden state $H_{i}^{A}$.
Figure 2: The DeepTalk model architecture.
Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better.
Model Seed-TTS $\downarrow$ $\downarrow$ $\downarrow$ Seed-TTS 1.12 2.25 7.59 CosyVoice 3.63 4.29 11.75 CosyVoice2 1.45 2.57 6.83 GLM-4-Voice 2.91 2.10 - VITA-1.5 8.44 2.63 - DeepTalk(topk=10) 1.87 7.34 7.53 DeepTalk(topk=1) 1.68 6.28 7.34 +DPO 1.41 3.25 7.29
Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite liu2024deepseek , as shown in Fig. 2. Each MoE layer is adaptively divided into modality expert groups for specialized training and learning of the corresponding modality. Specifically, the text expert group is responsible for generating written text such as mathematical code, while the speech expert group is responsible for generating spoken language such as emotional speech response. In addition, there is a shared modality expert group for learning common knowledge across both modalities such as daily conversations. The weights among modality expert groups are allocated through modality routing, as follows:
$P(X_{l}^{AT})_{i}=\frac{e^{f(X_{l}^{AT})}}{\sum_{j}^{M+N}e^{f(X_{l}^{AT})_{j}}},$ (5)
$MoE(X_{l}^{AT})=\sum_{i=1}^{L}P(X_{l}^{AT})_{i}\cdot e(X_{l}^{AT})_{i},$ (6)
where $X_{l}^{AT}$ denotes the averaged feature input of speech and text at the $l$-th layer, $L$ represents the total length of the sequence after concatenating speech and text, and $M$ and $N$ denote the number of speech experts and text experts, respectively. $f(x)=W\cdot x$ is the linear modality router to produce expert assignment probabilities and $W\in R^{h\times d}$. $h$ is the last dimension of hidden states and $d$ is the total number of experts.
Audio-Text Modeling Following parallel modeling paradigm defossez2024moshi ; xie2024mini ; chen2024slam ; gao2025lucy , we use different heads to process hidden states, generating both text and audio tokens chen2024slam ; defossez2024moshi in one decoding step. Since the input to the LLM is the average of the text and audio representations instead of native text representations, maintaining the original capabilities of the LLM can be challenging. To enhance models’ text ability, we apply batch parallel decoding xie2024mini ; gao2025lucy , Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response. The generated text tokens from the first sample are of better quality than those from the second and are used to substitute the latter in the second sample.
Audio Codec Decoder We use SNAC codec siuzdak2024snac to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz. Combining speech and text tokens results in eight layers of labels, so we use eight LM heads to simultaneously predict one text token and seven audio token at each decoding step. We follow the delay pattern of MusicGen (copet2024simple, ) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.
3.2 Adaptive Modality-Specific Mixture-of-Experts
The adaptive modality expert selection strategy is primarily designed for an original textual LLM, dynamically distinguishing between audio experts and text experts based on modality load. The goal is to ensure minimal impairment to the language capabilities of the original LLM. Specifically, as shown in Algorithm 1, an initial MoE model that has undergone stage 1 modality alignment training is used. For unimodal input data, the token load rates of all experts for both modalities are calculated, as illustrated in Figure 4. According to step 1 in Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts. As shown in Figure 4, for example, in the first layer, the expert with index 5 (starting from 0) is selected as an audio expert because its load rate for the audio modality is significantly higher than that for the text modality. This adaptive selection approach fully utilizes the text experts with low load in the original MoE LLM, while also ensuring that the selected experts are suitable for processing the audio modality. In this way, the strategy minimizes the impact on the original LLM while fitting the audio modality as much as possible.
Algorithm 1 Adaptive Modality Expert Partitioning and Training
0:  Data with audio input and audio-text output $D^{A}$, Data with text input and audio-text output $D^{T}$, The MoE LLM model $M$ after the first stage of modality alignment.
1:  Initialize $k$ = Number of audio experts, $N$ = Total number of experts.
2:  Step 1. Adaptive partitioning of modality experts
3:   Audio modality token load rate $E^{A}$ = $M(D^{A})$   # Inference with model $M$ on dataset $D^{A}$.
4:   Text modality token load rate $E^{T}$ = $M(D^{T})$   # Inference with model $M$ on dataset $D^{T}$.
5:   _$\_,AE$ = $topk(E^{A}*(1-E^{T}),k)$   # Compute the index $AE$ of the selected audio expert.
6:   _$\_,TE$ = $topk(E^{T}*(1-E^{A}),N-k)$   # Compute the index $TE$ of the selected text expert.
7:  Step 2. Expert specialization training for single modality
8:   $M[AE]$ = $M[AE](D^{A})$   # The audio expert undergoes specialized training on $D^{A}$.
9:   $M[TE]$ = $M[TE](D^{T})$   # The text expert undergoes specialized training on $D^{T}$.
10:  Stpe 3. Joint training of multi-modality experts
11:   $M^{\prime}$ = $M(D^{A}\cup D^{T})$
12:  return $M^{\prime}$, the MoE SLM model after modality expert partitioning and training.
3.3 Training Strategy for Modality Experts
In order to fully leverage the capabilities of both text and audio experts, it is necessary to conduct specialization training for the designated modality experts, enabling each group of modality experts to excel at processing inputs from their respective modalities. Finally, to handle multimodal information inputs simultaneously, cross-modal instruction data is used for joint training of the modality experts. The specific training process is illustrated in Figure 3.
Figure 3: The training process consists of three stages: stage 1 is for modality alignment, stage 2 is for training single modality experts, and stage 3 is for joint training of the modality experts.
Modality Aligment During the modality alignment stage, audio ASR data is used to align the semantic spaces of speech and text. In this stage, the main focus is on training the Adapter that connects the Audio Encoder and the LLM. Through modality alignment training, it is ensured that speech can be understood by the LLM in the same way as the meaning expressed by text.
Unimodal Expert Specialization Training The main purpose of unimodal specialization training is to enable modality experts to focus more on their own modality domains, so that they can fully utilize their expertise when collaborating with other multimodal experts. Since audio, like text, covers a wide range of domains, there are also multiple audio experts, each responsible for learning specialized knowledge in different areas, such as pitch, prosody, dialect, speech rate, and emotion. As shown in step 2 of Algorithm 1, audio and text experts are specialized and trained on their respective modalities. Since only unimodal specialization learning is required, the router is kept frozen to prevent it from becoming biased toward a single modality. During the specialization of audio experts, the router weight scores corresponding to text experts are set to zero. The remaining router weight scores are then normalized and redistributed among the audio experts. The same process is applied when specializing the text experts.
Joint Training of Modality Experts Multimodal joint training is mainly designed to enable the model to jointly train unimodal experts so that they can collaboratively process multimodal input information. As shown in Stage 3 of Figure 3, this stage uses cross-modal outputs from both modalities to jointly train the audio and text experts. Since all experts participate in training at this stage, the router is also unfrozen and trained, allowing it to better learn how to handle modality routing when processing multimodal joint inputs.
3.4 Audio Generation with Reinforcement Learning
To broaden the coverage of speakers and domains, it is inevitable that the pre-training data contains label noise and pronunciation errors, which can lead to hallucinations in the model. We observed that the quality of generated response audio is not consistent under different sampling parameters. To ensure more stable quality in response audio generation, we introduce a reinforcement learning stage to improve the stability of speech generation. Specifically, we use the text from the LibriSpeech panayotov2015librispeech set and have the model paraphrase it to obtain response audio and corresponding text. The response audio is then transcribed into recognized text using Whisper-large, and the word error rate (WER) is calculated as the reward score. All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) rafailov2023direct training as follows:
$\textit{L}_{DPO}(P_{\theta};P_{ref})=-\textit{E}_{(x,y_{w}^{A},y_{l}^{A})\sim D}[log\sigma(\beta log\frac{P_{\theta}(y_{w}^{A}|x))}{P_{ref}(y_{w}^{A}|x))}-\beta log\frac{P_{\theta}(y_{l}^{A}|x))}{P_{ref}(y_{l}^{A}|x))}],$ (7)
where $\theta$ is the policy model and $ref$ is the reference model. $x$ denotes the text input of the audio to be synthesized, $y_{w}^{A}$ denotes a good speech sequence, and $y_{l}^{A}$ denotes a bad speech sequence.
4 Experimental Results and Analysis
4.1 Experiment Settings
Models The overall architecture of the DeepTalk model is shown in Figure 2. We use the pre-trained DeepSeek-V2-Lite as the initial MoE-based LLM backbone, Whisper-medium as the audio encoder, and an MLP-based adapter to connect the audio encoder and the LLM. DeepTalk contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts. A sparse routing mechanism is adopted, where each token is processed by 6 experts. The model consists of 27 layers in total: the first layer is a dense layer, followed by 26 MoE layers. Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention dao2023flashattention with 16 attention heads. In stage 1, the learning rate is set to 2e-5. Since the original LLM has not undergone any pre-training on audio modalities, the learning rate is set to 1e-4 in stage 2.1 to accelerate convergence, 2e-5 in stage 2.2, and 5e-5 in stage 3. We use the AdamW loshchilov2017decoupled optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95. The final evaluation is conducted on LLM benchmarks, SQA, ASR, and TTS tasks.
Datasets In stage 1, WenetSpeech zhang2022wenetspeech is used for modality alignment. In stage 2.1, AudioQA-1M gao2025lucy is used for specialized training of audio experts. In stage 2.2, inhouse-60K, databricks-dolly-15k dolly2023introducing , MathInstruct yue2023mammoth (262K) and camel-ai-math li2023camel (50K) are used for specialized training of text experts. In stage 3, AudioQA-1M is used for joint training of modality experts. For the RL stage, we sampled 28K text entries from LibriSpeech panayotov2015librispeech to construct the audio preference pair dataset.
4.2 Evaluations
Evaluation on Text to Text
Figure 4: The modality load of the experts in the first layers.
The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/. As shown in Table 3, modular MLLMs cause minimal damage to the language abilities of the original text LLMs, primarily because the LLM is only responsible for generating text tokens. However, native MLLMs significantly impair the language capabilities of the original text LLMs, with an average relative loss of over 20% compared to their backbone LLMs. The DeepTalk model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
Evaluation on Speech to Text
Table 2: Performance on Spoken Question Answering
Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo chen2025minmo 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio huang2025step 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni wang2024freeze 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi defossez2024moshi 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice zeng2024glm 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY gao2025lucy 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend zhong2024moextend 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepTalk 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
The evaluation of speech-to-text is divided into two tasks: Spoken QA and ASR. As shown in the S $\rightarrow$ T column of Table 2, thanks to the preservation of the original LLM’s language capabilities by modular MLLMs, these models generally achieve leading performance. In contrast, native MLLMs cause significant degradation to the original LLM, resulting in increased hallucinations and thus lower scores. However, DeepTalk effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent. As a result, DeepTalk leads among native MLLMs. The ASR capabilities of the models are shown in Table 4, where DeepTalk, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
Table 3: Performance of Text to Text on LLM benchmark, * indicates that the average metric is calculated only on those test sets for which Qwen2.5-Omni has values. HS, WG, HEval denotes Hellaswag, Winogrande, HumanEval, respectively. “#AP” denotes number of activated parameters.
[…]
Table 4: Performance of Speech to Text on ASR. The metric for Chinese is CER, and for English WER. Lower values are better.
Model #A-Params WenetSpeech AIShell LibriSpeech
test-net test-meeting test test-clean test-other
Qwen2-Audio-base chu2024qwen2 7B 7.64 8.40 1.52 1.74 4.04
Baichuan-Audio-base li2025baichuan 7B 10.13 13.28 1.93 3.02 6.04
VITA fu2025vita 7B 12.15 16.53 - 8.14 18.41
Step-Audio-chat huang2025step 130B 9.47 10.83 2.14 3.19 10.67
Qwen2.5-Omni xu2025qwen2 7B 6.04 7.71 1.13 2.37 4.21
GLM-4-Voice zeng2024glm 9B - - 3.02 2.10 4.90
DeepTalk 2.4B 8.98 10.67 2.42 3.25 8.21
Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepTalk, where "topk" denotes the sampling parameter for audio tokens. Since the speech dialogue data used by DeepTalk contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English. After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in Recall@1, thereby making high-quality audio generation more stable. However, for more difficult tasks (test-hard), the supervised fine-tuning (SFT) model itself may not generate the correct result within the entire search space, so applying DPO brings little to no benefit.
Evaluation on Speech to Speech The evaluation of speech-to-speech is shown in the S $\rightarrow$ S column of Table 2. MoExtend refers to the results obtained by extending the original Deepseek-V2-Lite model with additional experts, adding six audio experts. However, since expanding the number of experts alters the distribution of the router, the performance is relatively suboptimal. DeepTalk, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts. This allows the audio experts to specialize while retaining the capabilities of the original text experts, thus achieving better performance among native MLLMs.
Latency We deployed DeepTalk on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module. The reported numbers are averaged over more than 10 samples. With a good network connection, the end-to-end latency of the entire system in half-duplex mode is 0.4362 seconds. The time cost for generating the first audio chunk is 0.3417 seconds. By subtracting the first chunk cost from the total latency, we estimate the network transmission cost to be 0.1062 seconds. The time cost for decoding a single token is 0.0205 seconds. All time measurements were conducted on a GPU with 80GB of memory.
5 Conclusion
To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepTalk, a MoE framework for adaptive modality expert learning. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs a performance loss lower than peer native MLLMs and on par with modular ones. Meanwhile, the end-to-end dialogue latency remains within 0.5 seconds, ensuring a seamless and intelligent speech interaction experience. To the best of our knowledge, this is the first MoE-based native multimodal large speech interaction model, which opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
Appendix A Dataset Details
As shown in Table 5, we provide the data and quantities used for training at each stage. Stage 1’s training data consists of speech-text pairs, which are used to align the semantic spaces of the speech and text modalities. AudioQA-1M contains conversational instruction data, with both input and output encompassing two modalities, aimed at fine-tuning the model to enhance its speech dialogue capabilities. Stage 2.2 utilizes pure text data to maintain the language proficiency of the text expert. During the RL stage, data from Librispeech is used to construct DPO preference pair data for speech generation.
Table 5: List of datasets used for each training stage.
Stage Dataset Size
1 WenetSpeech 10Kh
2.1 AudioQA-1M 1000K
2.2 In-House60k 60K
MathInstruct 262K
camel_ai_math 50K
databricks_dolly_15k 15K
3 AudioQA-1M 1000K
In-House60k 60K
MathInstruct 262K
camel_ai_math 50K
databricks_dolly_15k 15K
RL LibriSpeech-DPO 28K
Appendix B Experimental Setup Details
Table 6 outlines the main hyperparameters used during each stage of training. All ablation studies were conducted using a total batch size of 128, with learning rates of 1e-4, 2e-5, and 5e-5.
Table 6: Hyperparameters used in each stage training on DeepTalk.
Hyperparameter Stage 1 Stage 2.1 Stage 2.2 Stage 3 Stage RL
Learning rate 2e-5 1e-4 2e-5 5e-5 2e-5
LR schedule Cosine Cosine Cosine Cosine Cosine
Batchsize per GPU 16 8 8 8 8
Zero Zero2 Zero2-offload Zero2 Zero2-offload Zero3
Optimizer AdamW AdamW AdamW AdamW AdamW
Appendix C Ablation Study
C.1 Audio Experts Number
We conducted ablation experiments with varying numbers of audio experts, as shown in Table 7. The results demonstrate that as the number of audio experts increases, the language ability tends to be compromised, whereas the speech dialogue capability is enhanced. However, once the number of audio experts surpasses 24, both speech dialogue and language abilities begin to decline. This suggests that at this point, too many resources are allocated to audio experts at the expense of the text experts of the original text LLM, leading to a decrease in language ability. Moreover, the speech dialogue capability also falls after reaching a bottleneck concerning language ability.
Table 7: Ablation study on the impact of the number of audio experts on the capabilities of the two modalities.
Experts Num LLaMA Question Web Question TriviaQA Avg
T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S
6 69.0 59.7 46.6 23.1 58.2 27.5 57.9 36.8
12 61.0 60.0 39.2 26.7 49.7 29.1 50.0 38.6
18 59.0 59.0 32.1 28.3 42.6 33.7 44.6 40.3
24 48.0 48.0 23.7 22.9 31.9 30.2 34.5 33.7
C.2 The Modality Load of Different Layer Experts
Figure 5: The modality load of the experts in the middle layers.
Figure 6: The modality load of the experts in the last layers.
The expert load across different layers is illustrated in Figures 5 and 6, representing the modality load of an intermediate MoE layer and the final MoE layer, respectively. It is evident that there are significant differences in modality load across different layers. This indicates that the indices of audio experts in each layer are not the same.
C.3 Modality-Specific MoE
Table 8: Performance on Spoken Question Answering
Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
As shown in Table 8, we compared two settings: PureMoE without specifying modality experts and Random Modality-Specific MoE with randomly assigned experts. We found that if modality experts are not set, the model can only rely on self-learning the modality boundaries. During the early stages of training, the model does not clearly understand these boundaries, and the inclusion of audio data may significantly harm important text experts, leading to poorer final performance. The random assignment of modality experts was attempted three times, but results were inconsistent. If suitable experts were chosen as audio experts, the performance could be quite good, as seen in the second and third instances. However, if some crucial text experts were chosen as audio experts, the initial LLM suffers substantial damage, resulting in noticeable performance degradation. Using an adaptive modality assignment method effectively selects which experts are suitable as audio experts and text experts, achieving better performance.
Appendix D Border Impact
DeepTalk leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs. These limitations include the potential to generate inaccurate information or biased outputs. To address these issues, we enhanced the model’s speech interaction capabilities using AudioQA and performed speech-language instruction tuning on high-quality datasets. Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepTalk model in any downstream applications.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The contributions and scope of this paper have already been described in detail in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [N/A]
  9. Justification: No apparent limitations were identified in this study.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [No]
  14. Justification: This paper is application-oriented, and we have conducted extensive experiments to validate our conclusions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides a detailed description of the experimental procedure, and the experiments can be reproduced with the open-sourced code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide all the open-source datasets and codes that we used for training and evaluation.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all the training and inference details for reproducing and understanding our results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: For the end-to-end dialogue latency statistics of DeepTalk, we conducted 5 rounds of measurements, each averaging over 10 samples.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the details of the compute resources settings to reproduce our results.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We strictly follow the code of ethics during the project.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss societal impacts in the Appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We provide usage guidelines for users to adhere when accessing to our model.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We strictly follow the original licenses of existing assets including datasets and codes.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our models and codes are well documented for reproduction and protected by licenses.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not include crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not conduct any experiments with human subjects so this item is not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We use an LLM for grammar checking.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepTalk model in any downstream applications.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The contributions and scope of this paper have already been described in detail in the abstract and introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [N/A]
-  9. Justification: No apparent limitations were identified in this study.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [No]
-  14. Justification: This paper is application-oriented, and we have conducted extensive experiments to validate our conclusions.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The paper provides a detailed description of the experimental procedure, and the experiments can be reproduced with the open-sourced code.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We provide all the open-source datasets and codes that we used for training and evaluation.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We provide all the training and inference details for reproducing and understanding our results.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: For the end-to-end dialogue latency statistics of DeepTalk, we conducted 5 rounds of measurements, each averaging over 10 samples.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide the details of the compute resources settings to reproduce our results.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We strictly follow the code of ethics during the project.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We discuss societal impacts in the Appendix.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: We provide usage guidelines for users to adhere when accessing to our model.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We strictly follow the original licenses of existing assets including datasets and codes.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: Our models and codes are well documented for reproduction and protected by licenses.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Our paper does not include crowdsourcing or human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not conduct any experiments with human subjects so this item is not applicable.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: We use an LLM for grammar checking.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
-DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE
+DeepOmni: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE
 Abstract
@@
 However, native MLLMs suffer from catastrophic forgetting and performance degradation because the available paired speech-text data is insufficient to support the pretraining of MLLMs compared to the vast amount of text data required to pretrain text LLMs.
-To address this issue, we propose DeepTalk, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture.
-DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM.
+To address this issue, we propose DeepOmni, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture.
+DeepOmni first adaptively distinguishes modality experts according to their modality load within the LLM.
 Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training.
-As a result, DeepTalk incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs.
+As a result, DeepOmni incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs.
 Meanwhile, the end-to-end dialogue latency remains within $0.5$ seconds, ensuring a seamless and intelligent speech interaction experience.
-Code and models are released at https://github.com/talkking/DeepTalk.
 1 Introduction
-Since GPT-4o GPT4o has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models.
-As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) radford2023robust ; bai2024seed ; povey2011kaldi , the resulting text is then comprehended by a large language model (LLM) liu2024deepseek ; yang2024qwen2 ; glm2024chatglm to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) du2024cosyvoice ; ren2019fastspeech ; wang2025spark model.
+Since GPT-4o OpenAI (2024) has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models.
+As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) Radford et al.
+(2023); Bai et al.
+(2024); Povey et al.
+(2011), the resulting text is then comprehended by a large language model (LLM) Liu et al.
+(2024); Yang et al.
+(2024); GLM et al.
+(2024) to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) Du et al.
+(2024); Ren et al.
+(2019); Wang et al.
+(2025) model.
 However, this cascaded structure results in high latency, and errors from each module can accumulate and propagate, leading to degraded performance.
@@
 At the same time, this approach significantly reduces interaction latency and provides a more natural and smooth voice interaction experience.
-Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal chen2025minmo .
-Representatives of the former include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , Freeze-Omni wang2024freeze , and VITA-1.5 fu2025vita .
+Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal Chen et al.
+(2025).
+Representatives of the former include Qwen2.5-Omni Xu et al.
+(2025), Minmo Chen et al.
+(2025), LLaMA-Omni Fang et al.
+(2024), Freeze-Omni Wang et al.
+(2024), and VITA-1.5 Fu et al.
+(2025).
 These models connect a speech encoder and a speech decoder to the LLM to handle audio understanding and audio generation tasks separately, placing semantic understanding within the LLM module.
@@
 Furthermore, since the LLM itself does not require retraining, much less speech-text paired pretraining data is needed.
-Representative of the latter including Mini-Omni xie2024mini , GLM-4-Voice zeng2024glm , Moshi defossez2024moshi , and LUCY gao2025lucy use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens.
+Representative of the latter including Mini-Omni Xie & Wu (2024), GLM-4-Voice Zeng et al.
+(2024), Moshi Défossez et al.
+(2024), LUCY Gao et al.
+(2025) and Step-Audio2 Wu et al.
+(2025) use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens.
 Compared to modular aligned multimodal approaches, native ones greatly reduces the latency of speech token output.
@@
 Therefore, in real-world scenarios where audio data is much less abundant than text data, native multimodal models are prone to catastrophic forgetting, severely impairing the original LLM’s language capabilities.
-In this paper, to alleviate the catastrophic forgetting faced by native multimodal models zhai2023investigating , we propose DeepTalk, a model that applies the mixture of experts (MoE) for multimodal learning li2025uni ; li2024cumo ; zhong2024moextend ; shen2024mome to effectively isolate modality-specific knowledge.
+In this paper, to alleviate the catastrophic forgetting faced by native multimodal models Zhai et al.
+(2023), we propose DeepOmni, a model that applies the mixture of experts (MoE) for multimodal learning Li et al.
+(2025b; 2024); Zhong et al.
+(2024); Shen et al.
+(2024) to effectively isolate modality-specific knowledge.
 Since most speech dialogue data are conversational and colloquial, while text LLMs are primarily trained on formal written text, training LLMs with colloquial speech dialogue data can lead to less standardized written outputs, resulting in decreased language performance on certain LLM benchmarks.
@@
 Experts with a high speech token load but a low text token load are selected as audio experts.
-After several iterations of modality expert selection and training, DeepTalk achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
+After several iterations of modality expert selection and training, DeepOmni achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
 In summary, the contributions of this paper are as follows:
-  • We propose DeepTalk, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge.
+  • We propose DeepOmni, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge.
 At the same time, modalities can collaborate to jointly output multimodal results, enabling both colloquial and formal language to coexist within a single model.
@@
 It opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
-Figure 1: Comparison of DeepTalk and existing voice interaction systems.
-Compared to modular MLLMs, DeepTalk embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM.
-Through adaptive modality expert selection and a three-stage training process for modality experts, DeepTalk effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
+Figure 1: Comparison of DeepOmni and existing voice interaction systems.
+Compared to modular MLLMs, DeepOmni embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM.
+Through adaptive modality expert selection and a three-stage training process for modality experts, DeepOmni effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
 2 Related Work
@@
 Such systems not only suffer from high interaction latency, but also accumulate errors from each module, leading to a less satisfactory overall user experience.
-Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach ji2024wavchat .
-Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM kong2020panns ; chu2024qwen2 ; chu2023qwen ; das2024speechverse .
+Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach Ji et al.
+(2024a).
+Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM Kong et al.
+(2020); Chu et al.
+(2024; 2023); Das et al.
+(2024).
 This approach not only provides the LLM with richer speech information than using pure text input, but also reduces the latency caused by ASR recognition.
 However, these models still require an additional TTS module to achieve end-to-end speech interaction.
-To provide rich speech information, LLaMA-Omni fang2024llama feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens.
-Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss graves2006connectionist is used to align the upsampled hidden states with the audio token outputs.
+To provide rich speech information, LLaMA-Omni Fang et al.
+(2024) feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens.
+Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss Graves et al.
+(2006) is used to align the upsampled hidden states with the audio token outputs.
 Meanwhile, to reduce latency for real-time speech interaction, a predefined chunk size is set to enable the vocoder to synthesize speech in a streaming manner.
-Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of $82$Hz and adopts a delayed decoding strategy copet2023simple ; peng2024voicecraft ; lyth2024natural to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference.
-Minmo chen2025minmo integrates a language model with an autoregressive CosyVoice2 decoder du2024cosyvoice .
+Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of 8282Hz and adopts a delayed decoding strategy Copet et al.
+(2023); Peng et al.
+(2024); Lyth & King (2024) to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference.
+Minmo Chen et al.
+(2025) integrates a language model with an autoregressive CosyVoice2 decoder Du et al.
+(2024).
 The hidden states of the language model are fed into the CosyVoice2 decoder to autoregressively predict speech tokens, which are then passed to a vocoder for waveform synthesis.
 Fully end-to-end approaches integrate the TTS module into the LLM, allowing the LLM itself to generate both text and speech tokens, rather than relying on a separate speech decoder or TTS module for speech synthesis.
-Discrete speech tokens are typically generated by neural audio codecs defossez2022high ; ji2024wavtokenizer ; zhang2023speechtokenizer or self-supervised models lakhotia2021generative ; hsu2021hubert ; chen2022wavlm .
+Discrete speech tokens are typically generated by neural audio codecs Défossez et al.
+(2022); Ji et al.
+(2024b); Zhang et al.
+(2023) or self-supervised models Lakhotia et al.
+(2021); Hsu et al.
+(2021); Chen et al.
+(2022).
 Currently, there are two mainstream modeling paradigm for end-to-end speech systems.
-The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice zeng2024glm and Spirit-LM nguyen2025spirit , where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens.
-The second is parallel audio-text modeling, represented by models such as Mini-Omni xie2024mini , Moshi defossez2024moshi , Slam-Omni chen2024slam and LUCY gao2025lucy .
+The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice Zeng et al.
+(2024), Baichuan-Audio Li et al.
+(2025a) and Spirit-LM Nguyen et al.
+(2025), where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens.
+The second is parallel audio-text modeling, represented by models such as Mini-Omni Xie & Wu (2024), Moshi Défossez et al.
+(2024), Slam-Omni Chen et al.
+(2024) and LUCY Gao et al.
+(2025).
 In this paradigm, text and audio tokens can be generated in parallel: while generating text tokens, multiple LM heads are used to simultaneously predict multiple audio tokens.
@@
 Modular Speech Language Models  Modular large speech models are primarily built on an architecture where a LLM is directly connected to an audio encoder and decoder through adapters.
-Examples include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , and Freeze-Omni wang2024freeze , as shown in Fig.
+Examples include Qwen2.5-Omni Xu et al.
+(2025), Minmo Chen et al.
+(2025), LLaMA-Omni Fang et al.
+(2024), and Freeze-Omni Wang et al.
+(2024), as shown in Fig.
 1 b).
@@
 As a result, the language capabilities of the LLM are largely preserved.
-Additionally, Minmo employs Low-Rank Adaptation (LoRA) hu2022lora for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities.
+Additionally, Minmo employs Low-Rank Adaptation (LoRA) Hu et al.
+(2022) for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities.
 However, such modular speech-language models face several challenges.
@@
 This means that the expression of paralinguistic information in speech is completely dependent on the speech decoder, resulting in underutilization of the LLM.
-Furthermore, there are additional issues such as low deployment efficiency and the complexity of balancing the capacities of the LLM and the SLM.
+Furthermore, there are additional issues such as low deployment efficiency.
+In modular speech language models paradigm, the LLM is trained to model only text distribution, or formally $P(T_{out}|T_{in}\cup S_{in})$, and the audio decoder is trained to model $P(S_{out}|T_{out})$.
 Native Speech Language Models Native speech language models reconstruct the LLM into an SLM, enabling the LLM to output both text and speech tokens simultaneously, as shown in Fig.
 1 c).
-Currently, most mainstream native multimodal approaches such as Mini-Omni xie2024mini , Moshi defossez2024moshi , GLM-4-Voice zeng2024glm , and LUCY gao2025lucy have the LLM handle both text and speech decoding.
+Currently, most mainstream native multimodal approaches such as Mini-Omni Xie & Wu (2024), Moshi Défossez et al.
+(2024), GLM-4-Voice Zeng et al.
+(2024), and LUCY Gao et al.
+(2025) have the LLM handle both text and speech decoding.
 This approach offers low latency, easy deployment, and allows a unified model to process both audio and text decoding, which better aligns with end-to-end requirements.
 However, without large-scale speech-text paired data, retraining the LLM in this way can lead to catastrophic forgetting and performance degradation.
+In native speech language models paradigm, the LLM is trained to model $P(T_{out}\cup S_{out}|T_{in}\cup S_{in})$ .
+Clearly in the native paradigm, the speech-text joint distribution space for the LLM to learn is more complex than text distribution.
+It is also for this reason in modular paradigm the LLM does not need to be restructed to a SLM and its language capacity is largely preserved.
 2.2 Multimodal Mixture-of-Experts
-BEiT-3 wang2023image , VLMo bao2022vlmo , and Uni-MoE li2025uni employ specific modality expert groups to capture modality-specific information.
-MoExtend zhong2024moextend further expands modality experts horizontally on top of existing LLMs to capture additional modality information.
-CuMo li2024cumo extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME shen2024mome uses multiple feature encoders as modality expert groups to encode multi-dimensional features.
-VL-MoE shen2023scaling and MoE-LLaVa lin2024moe introduce mixture-of-experts (MoE) to improve training and deployment.
-MoMa lin2024moma pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens.
+Multimodal Mixture-of-Experts refers to using a MoE model to learn multimodal knowledge.
+BEiT-3 Wang et al.
+(2023), VLMo Bao et al.
+(2022), and Uni-MoE Li et al.
+(2025b) employ specific modality expert groups to capture modality-specific information.
+MoExtend Zhong et al.
+(2024) further expands modality experts horizontally on top of existing LLMs to capture additional modality information.
+CuMo Li et al.
+(2024) extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME Shen et al.
+(2024) uses multiple feature encoders as modality expert groups to encode multi-dimensional features.
+VL-MoE Shen et al.
+(2023) and MoE-LLaVa Lin et al.
+(2024a) introduce mixture-of-experts (MoE) to improve training and deployment.
+MoMa Lin et al.
+(2024b) pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens.
 Inspired by these works, we introduce multimodal mixture-of-experts (speech experts and text experts) into end-to-end large speech interaction models for native MLLM pretraining.
@@
 The model’s text and speech responses are represented as $Y^{T}\in V$ and $Y^{A}\in U$, where $V$ and $U$ denote the text vocabulary and the speech codec vocabulary, respectively.
-The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model, as shown below:
-$H_{i}^{A}=Adapter(Encoder(X_{i}^{A})),$ (1)
-$H_{i}^{T}=Embedding(X_{i}^{T}),$ (2)
-$H_{i}=mean(H_{i}^{A},H_{i}^{T}),$ (3)
-where $mean$ here refers to taking the average of the two embeddings. The loss L of the model over N samples can be defined as:
+The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model.
+The loss L of the model over N samples can be defined as:
 $\textit{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}logP(Y_{i,t}^{T},Y_{i,t}^{A}|Y_{i,<t}^{T},Y_{i,<t}^{A},H_{i}),$ (1)
-where $T_{i}$ donates the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
-Audio Encoder and Adapter We use Whisper-medium radford2023robust as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads.
+where $T_{i}$ denotes the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
+Audio Encoder and Adapter We use Whisper-medium Radford et al.
+(2023) as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads.
 It includes multiple convolutional downsampling layers that downsample the speech features by a factor of 4.
 We use an 80-dimensional log Mel-filterbank with a 25ms window length computed every 10ms as the input to the audio encoder.
-A convolution-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features.
-Finally, for the input $X_{i}^{A}$, as shown in Equation 1, we obtain the hidden state $H_{i}^{A}$.
-Figure 2: The DeepTalk model architecture.
-Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better.
-Model Seed-TTS $\downarrow$ $\downarrow$ $\downarrow$ Seed-TTS 1.12 2.25 7.59 CosyVoice 3.63 4.29 11.75 CosyVoice2 1.45 2.57 6.83 GLM-4-Voice 2.91 2.10 - VITA-1.5 8.44 2.63 - DeepTalk(topk=10) 1.87 7.34 7.53 DeepTalk(topk=1) 1.68 6.28 7.34 +DPO 1.41 3.25 7.29
-Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite liu2024deepseek , as shown in Fig.
+A MLP-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features.
+Finally, for the input $X_{i}^{A}$, we obtain the hidden state $H_{i}^{A}=Adapter(Encoder(X_{i}^{A}))$.
+Figure 2: The DeepOmni model architecture.
+Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better. The metric for Chinese is character error rate (CER), and the metric for English is word error rate (WER).
+[…]
+Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite Liu et al.
+(2024), as shown in Fig.
 2.
@@
 The weights among modality expert groups are allocated through modality routing, as follows:
-$P(X_{l}^{AT})_{i}=\frac{e^{f(X_{l}^{AT})}}{\sum_{j}^{M+N}e^{f(X_{l}^{AT})_{j}}},$ (5)
-$MoE(X_{l}^{AT})=\sum_{i=1}^{L}P(X_{l}^{AT})_{i}\cdot e(X_{l}^{AT})_{i},$ (6)
-where $X_{l}^{AT}$ denotes the averaged feature input of speech and text at the $l$-th layer, $L$ represents the total length of the sequence after concatenating speech and text, and $M$ and $N$ denote the number of speech experts and text experts, respectively.
+$P(h_{l})_{i}=\frac{e^{f(h_{l})_{i}}}{\sum_{j}^{M}e^{f(h_{l})_{i,j}}},$ (2)
+$MoE(h_{l})=\sum_{i=1}^{T}P(h_{l})_{i}\cdot e(h_{l})_{i},$ (3)
+where $h_{l}$ denotes the averaged feature input of speech and text at the $l$-th layer, $T$ represents the total length of the sequence after concatenating speech and text, and $M$ denotes the total number of speech and text experts combined, respectively.
 $f(x)=W\cdot x$ is the linear modality router to produce expert assignment probabilities and $W\in R^{h\times d}$.
 $h$ is the last dimension of hidden states and $d$ is the total number of experts.
-Audio-Text Modeling Following parallel modeling paradigm defossez2024moshi ; xie2024mini ; chen2024slam ; gao2025lucy , we use different heads to process hidden states, generating both text and audio tokens chen2024slam ; defossez2024moshi in one decoding step.
+Audio-Text Modeling Following parallel modeling paradigm Défossez et al.
+(2024); Xie & Wu (2024); Chen et al.
+(2024); Gao et al.
+(2025), we use different heads to process hidden states, generating both text and audio tokens Chen et al.
+(2024); Défossez et al.
+(2024) in one decoding step.
 Since the input to the LLM is the average of the text and audio representations instead of native text representations, maintaining the original capabilities of the LLM can be challenging.
-To enhance models’ text ability, we apply batch parallel decoding xie2024mini ; gao2025lucy , Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response.
+To enhance models’ text ability, we apply batch parallel decoding Xie & Wu (2024); Gao et al.
+(2025), Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response.
 The generated text tokens from the first sample are of better quality than those from the second and are used to substitute the latter in the second sample.
-Audio Codec Decoder We use SNAC codec siuzdak2024snac to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz.
+Audio Codec Decoder We use SNAC codec Siuzdak et al.
+(2024) to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz.
 Combining speech and text tokens results in eight layers of labels, so we use eight LM heads to simultaneously predict one text token and seven audio token at each decoding step.
-We follow the delay pattern of MusicGen (copet2024simple, ) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.
+We follow the delay pattern of MusicGen (Copet et al., 2024) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.
 3.2 Adaptive Modality-Specific Mixture-of-Experts
@@
 Specifically, as shown in Algorithm 1, an initial MoE model that has undergone stage 1 modality alignment training is used.
-For unimodal input data, the token load rates of all experts for both modalities are calculated, as illustrated in Figure 4.
-According to step 1 in Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts.
-As shown in Figure 4, for example, in the first layer, the expert with index 5 (starting from 0) is selected as an audio expert because its load rate for the audio modality is significantly higher than that for the text modality.
+For unimodal input data, the token load rates of all experts for both modalities are calculated.
+According to Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts.
 This adaptive selection approach fully utilizes the text experts with low load in the original MoE LLM, while also ensuring that the selected experts are suitable for processing the audio modality.
 In this way, the strategy minimizes the impact on the original LLM while fitting the audio modality as much as possible.
-Algorithm 1 Adaptive Modality Expert Partitioning and Training
-0:  Data with audio input and audio-text output $D^{A}$, Data with text input and audio-text output $D^{T}$, The MoE LLM model $M$ after the first stage of modality alignment.
-1:  Initialize $k$ = Number of audio experts, $N$ = Total number of experts.
-2:  Step 1. Adaptive partitioning of modality experts
-3:   Audio modality token load rate $E^{A}$ = $M(D^{A})$   # Inference with model $M$ on dataset $D^{A}$.
-4:   Text modality token load rate $E^{T}$ = $M(D^{T})$   # Inference with model $M$ on dataset $D^{T}$.
-5:   _$\_,AE$ = $topk(E^{A}*(1-E^{T}),k)$   # Compute the index $AE$ of the selected audio expert.
-6:   _$\_,TE$ = $topk(E^{T}*(1-E^{A}),N-k)$   # Compute the index $TE$ of the selected text expert.
-7:  Step 2. Expert specialization training for single modality
-8:   $M[AE]$ = $M[AE](D^{A})$   # The audio expert undergoes specialized training on $D^{A}$.
-9:   $M[TE]$ = $M[TE](D^{T})$   # The text expert undergoes specialized training on $D^{T}$.
-10:  Stpe 3. Joint training of multi-modality experts
-11:   $M^{\prime}$ = $M(D^{A}\cup D^{T})$
-12:  return $M^{\prime}$, the MoE SLM model after modality expert partitioning and training.
+Algorithm 1 Adaptive Modality Expert Partitioning
+1:
+  • Multimodal dataset with audio input: $\mathcal{D}^{A}=\{(x_{i}^{A},y_{i}^{A,T})\}_{i=1}^{n}$
+  • Multimodal dataset with text input: $\mathcal{D}^{T}=\{(x_{j}^{T},y_{j}^{A,T})\}_{j=1}^{m}$
+  • Pre-trained MoE-LLM $\pi_{\theta}$ after modality alignment
+2:Specialized expert allocation strategy $\{\mathbf{E}_{l}^{A},\mathbf{E}_{l}^{T}\}_{l=1}^{L}$ {$L$ is the total number of layers.}
+3:Initialize: $k\leftarrow\text{\# audio experts}$, $M\leftarrow\text{\# total experts}$, $C\leftarrow\text{\# activated experts}$
+4:Phase 1: Expert Selection Statistics
+5:Extract hidden states: $\mathbf{h}_{l}^{A}=\pi_{\theta}(\mathcal{D}^{A})$, $\mathbf{h}_{l}^{T}=\pi_{\theta}(\mathcal{D}^{T})$ {Layer-wise representations}
+[…]
+7:Initialize counting matrices: $\mathbf{E}^{A}_{l}\leftarrow\mathbf{0}_{M}$, $\mathbf{E}^{T}_{l}\leftarrow\mathbf{0}_{M}$ {Per-layer expert counters}
+8:for $(modality,\mathbf{h}_{l})$ in $\{(A,\mathbf{h}_{l}^{A}),(T,\mathbf{h}_{l}^{T})\}$ do
+9:for $i\leftarrow 1\text{ to }T^{modality}$ do
+10:  $\mathcal{E}_{l,i}^{modality}\leftarrow\mathrm{top\text{-}k}\big(\mathrm{MoE}(\mathbf{h}_{l}^{modality})_{i},C\big)$ {MoE using equation 3}
+11:  for $j\leftarrow 1\text{ to }\text{len}(\mathcal{E}_{l,i}^{modality})$ do
+12:   $\mathbf{E}^{modality}_{l}[\mathcal{E}_{l,i}^{modality}[j]]\mathrel{+}=1$ {Aggregate selections}
+13:  end for
+14:end for
+15:end for
+16:Phase 2: Modality Token Load Ratio Computation
+17:for $modality\in\{A,T\}$ do
+18:for $j\leftarrow 1\text{ to }M$ do
+19:  $\rho_{l,j}^{modality}\leftarrow\frac{\mathbf{E}^{modality}_{l,j}}{C\cdot T^{modality}}$ {Token load ratio $\rho\in[0,1]$}
+20:end for
+21:end for
+22:Phase 3: Partitioning Modality Experts Based on Modality Token Load
+23:for $j\leftarrow 1\text{ to }M$ do
+[…]
+[…]
+26:end for
+27:return $\pi_{\theta}(\mathrm{Audio\_Experts}_{l},\mathrm{Text\_Experts}_{l})$
 3.3 Training Strategy for Modality Experts
@@
 Through modality alignment training, it is ensured that speech can be understood by the LLM in the same way as the meaning expressed by text.
+The specific approach is to use a cross-entropy loss to align the audio and text modalities.
+Because the audio sequence for the same content is significantly longer than the text, in addition to downsampling the audio, we also use text padding tokens to align their lengths.
 Unimodal Expert Specialization Training The main purpose of unimodal specialization training is to enable modality experts to focus more on their own modality domains, so that they can fully utilize their expertise when collaborating with other multimodal experts.
@@
 To ensure more stable quality in response audio generation, we introduce a reinforcement learning stage to improve the stability of speech generation.
-Specifically, we use the text from the LibriSpeech panayotov2015librispeech set and have the model paraphrase it to obtain response audio and corresponding text.
+Specifically, we use the text from the LibriSpeech Panayotov et al.
+(2015) set and have the model paraphrase it to obtain response audio and corresponding text.
 The response audio is then transcribed into recognized text using Whisper-large, and the word error rate (WER) is calculated as the reward score.
-All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) rafailov2023direct training as follows:
+All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) Rafailov et al.
+(2023) training as follows:
 $\textit{L}_{DPO}(P_{\theta};P_{ref})=-\textit{E}_{(x,y_{w}^{A},y_{l}^{A})\sim D}[log\sigma(\beta log\frac{P_{\theta}(y_{w}^{A}|x))}{P_{ref}(y_{w}^{A}|x))}-\beta log\frac{P_{\theta}(y_{l}^{A}|x))}{P_{ref}(y_{l}^{A}|x))}],$ (4)
-where $\theta$ is the policy model and $ref$ is the reference model.
+where $\theta$ is the policy model and refref is the reference model.
 $x$ denotes the text input of the audio to be synthesized, $y_{w}^{A}$ denotes a good speech sequence, and $y_{l}^{A}$ denotes a bad speech sequence.
@@
 4.1 Experiment Settings
-Models The overall architecture of the DeepTalk model is shown in Figure 2.
+Models The overall architecture of the DeepOmni model is shown in Figure 2.
 We use the pre-trained DeepSeek-V2-Lite as the initial MoE-based LLM backbone, Whisper-medium as the audio encoder, and an MLP-based adapter to connect the audio encoder and the LLM.
-DeepTalk contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts.
+DeepOmni contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts.
 A sparse routing mechanism is adopted, where each token is processed by 6 experts.
 The model consists of 27 layers in total: the first layer is a dense layer, followed by 26 MoE layers.
-Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention dao2023flashattention with 16 attention heads.
+Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention Dao (2023) with 16 attention heads.
 In stage 1, the learning rate is set to 2e-5.
 Since the original LLM has not undergone any pre-training on audio modalities, the learning rate is set to 1e-4 in stage 2.1 to accelerate convergence, 2e-5 in stage 2.2, and 5e-5 in stage 3.
-We use the AdamW loshchilov2017decoupled optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95.
+We use the AdamW Loshchilov & Hutter (2017) optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95.
 The final evaluation is conducted on LLM benchmarks, SQA, ASR, and TTS tasks.
-Datasets In stage 1, WenetSpeech zhang2022wenetspeech is used for modality alignment.
-In stage 2.1, AudioQA-1M gao2025lucy is used for specialized training of audio experts.
-In stage 2.2, inhouse-60K, databricks-dolly-15k dolly2023introducing , MathInstruct yue2023mammoth (262K) and camel-ai-math li2023camel (50K) are used for specialized training of text experts.
+Datasets In stage 1, WenetSpeech Zhang et al.
+(2022) is used for modality alignment.
+In stage 2.1, AudioQA-1M Gao et al.
+(2025) is used for specialized training of audio experts.
+In stage 2.2, databricks-dolly-15k Dolly (2023), MathInstruct Yue et al.
+(2023) (262K) and camel-ai-math Li et al.
+(2023) (50K) are used for specialized training of text experts.
 In stage 3, AudioQA-1M is used for joint training of modality experts.
-For the RL stage, we sampled 28K text entries from LibriSpeech panayotov2015librispeech to construct the audio preference pair dataset.
+For the RL stage, we sampled 28K text entries from LibriSpeech Panayotov et al.
+(2015) to construct the audio preference pair dataset.
 4.2 Evaluations
-Evaluation on Text to Text
-Figure 4: The modality load of the experts in the first layers.
-The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/.
+Evaluation on Text to Text The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/.
 As shown in Table 3, modular MLLMs cause minimal damage to the language abilities of the original text LLMs, primarily because the LLM is only responsible for generating text tokens.
 However, native MLLMs significantly impair the language capabilities of the original text LLMs, with an average relative loss of over 20% compared to their backbone LLMs.
-The DeepTalk model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
+The DeepOmni model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
 Evaluation on Speech to Text
 Table 2: Performance on Spoken Question Answering
-Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo chen2025minmo 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio huang2025step 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni wang2024freeze 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi defossez2024moshi 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice zeng2024glm 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY gao2025lucy 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend zhong2024moextend 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepTalk 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
+Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo Chen et al.
+(2025) 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio Huang et al.
+(2025) 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni Wang et al.
+(2024) 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi Défossez et al.
+(2024) 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice Zeng et al.
+(2024) 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY Gao et al.
+(2025) 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend Zhong et al.
+(2024) 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepOmni 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
 The evaluation of speech-to-text is divided into two tasks: Spoken QA and ASR.
@@
 In contrast, native MLLMs cause significant degradation to the original LLM, resulting in increased hallucinations and thus lower scores.
-However, DeepTalk effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent.
-As a result, DeepTalk leads among native MLLMs.
-The ASR capabilities of the models are shown in Table 4, where DeepTalk, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
+However, DeepOmni effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent.
+As a result, DeepOmni leads among native MLLMs.
+The ASR capabilities of the models are shown in Table 4, where DeepOmni, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
 Table 3: Performance of Text to Text on LLM benchmark, * indicates that the average metric is calculated only on those test sets for which Qwen2.5-Omni has values.
@@
 Table 4: Performance of Speech to Text on ASR. The metric for Chinese is CER, and for English WER. Lower values are better.
-Model #A-Params WenetSpeech AIShell LibriSpeech
-test-net test-meeting test test-clean test-other
-Qwen2-Audio-base chu2024qwen2 7B 7.64 8.40 1.52 1.74 4.04
-Baichuan-Audio-base li2025baichuan 7B 10.13 13.28 1.93 3.02 6.04
-VITA fu2025vita 7B 12.15 16.53 - 8.14 18.41
-Step-Audio-chat huang2025step 130B 9.47 10.83 2.14 3.19 10.67
-Qwen2.5-Omni xu2025qwen2 7B 6.04 7.71 1.13 2.37 4.21
-GLM-4-Voice zeng2024glm 9B - - 3.02 2.10 4.90
-DeepTalk 2.4B 8.98 10.67 2.42 3.25 8.21
-Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepTalk, where "topk" denotes the sampling parameter for audio tokens.
-Since the speech dialogue data used by DeepTalk contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English.
-After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in Recall@1, thereby making high-quality audio generation more stable.
+Model #A-Params WenetSpeech AIShell LibriSpeech test-net test-meeting test test-clean test-other Qwen2-Audio-base Chu et al.
+(2024) 7B 7.64 8.40 1.52 1.74 4.04 Baichuan-Audio-base Li et al.
+(2025a) 7B 10.13 13.28 1.93 3.02 6.04 VITA Fu et al.
+(2025) 7B 12.15 16.53 - 8.14 18.41 Step-Audio-chat Huang et al.
+(2025) 130B 9.47 10.83 2.14 3.19 10.67 Qwen2.5-Omni Xu et al.
+(2025) 7B 6.04 7.71 1.13 2.37 4.21 GLM-4-Voice Zeng et al.
+(2024) 9B - - 3.02 2.10 4.90 DeepOmni 2.4B 8.98 10.67 2.42 3.25 8.21
+Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepOmni, where ”topk” denotes the sampling parameter for audio tokens.
+Since the speech dialogue data used by DeepOmni contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English.
+After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in pass@1, thereby making high-quality audio generation more stable.
 However, for more difficult tasks (test-hard), the supervised fine-tuning (SFT) model itself may not generate the correct result within the entire search space, so applying DPO brings little to no benefit.
@@
 However, since expanding the number of experts alters the distribution of the router, the performance is relatively suboptimal.
-DeepTalk, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts.
+DeepOmni, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts.
 This allows the audio experts to specialize while retaining the capabilities of the original text experts, thus achieving better performance among native MLLMs.
-Latency We deployed DeepTalk on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module.
+Latency We deployed DeepOmni on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module.
 The reported numbers are averaged over more than 10 samples.
@@
 5 Conclusion
-To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepTalk, a MoE framework for adaptive modality expert learning.
-DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM.
+To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepOmni, a MoE framework for adaptive modality expert learning.
+DeepOmni first adaptively distinguishes modality experts according to their modality load within the LLM.
 Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training.
-As a result, DeepTalk incurs a performance loss lower than peer native MLLMs and on par with modular ones.
+As a result, DeepOmni incurs a performance loss lower than peer native MLLMs and on par with modular ones.
 Meanwhile, the end-to-end dialogue latency remains within 0.5 seconds, ensuring a seamless and intelligent speech interaction experience.
 To the best of our knowledge, this is the first MoE-based native multimodal large speech interaction model, which opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
-Appendix A Dataset Details
+6 Reproducibility statement
+We have released all reproducible code in the supplementary materials and anonymized it.
+The details of the training data used and the hyperparameters are provided in Appendices A.1 and B, respectively.
+Appendix A Appendix
+A.1 Dataset Details
 As shown in Table 5, we provide the data and quantities used for training at each stage.
@@
 2.1 AudioQA-1M 1000K
-2.2 In-House60k 60K
-MathInstruct 262K
+2.2 MathInstruct 262K
 camel_ai_math 50K
@@
 3 AudioQA-1M 1000K
-In-House60k 60K
 MathInstruct 262K
@@
 Table 6 outlines the main hyperparameters used during each stage of training. All ablation studies were conducted using a total batch size of 128, with learning rates of 1e-4, 2e-5, and 5e-5.
-Table 6: Hyperparameters used in each stage training on DeepTalk.
+Table 6: Hyperparameters used in each stage training on DeepOmni.
 Hyperparameter Stage 1 Stage 2.1 Stage 2.2 Stage 3 Stage RL
@@
 Optimizer AdamW AdamW AdamW AdamW AdamW
-Appendix C Ablation Study
-C.1 Audio Experts Number
+B.1 Ablation Study
+B.1.1 Audio Experts Number
 We conducted ablation experiments with varying numbers of audio experts, as shown in Table 7.
@@
 24 48.0 48.0 23.7 22.9 31.9 30.2 34.5 33.7
-C.2 The Modality Load of Different Layer Experts
+B.1.2 The Modality Load of Different Layer Experts
+Figure 4: The modality load of the experts in the first layers.
 Figure 5: The modality load of the experts in the middle layers.
 Figure 6: The modality load of the experts in the last layers.
-The expert load across different layers is illustrated in Figures 5 and 6, representing the modality load of an intermediate MoE layer and the final MoE layer, respectively.
-It is evident that there are significant differences in modality load across different layers.
-This indicates that the indices of audio experts in each layer are not the same.
-C.3 Modality-Specific MoE
+The expert loads at different layers are shown in Figures  4,  5, and  6, corresponding to the modality loads of the first, middle, and last MoE layers, respectively.
+We can see that the modality loads differ substantially across layers, indicating that the indices of the audio experts vary from layer to layer.
+Moreover, the modality token loads at each layer also reveal the specialization of the modality-specific experts.
+B.1.3 Modality-Specific MoE
 Table 8: Performance on Spoken Question Answering
-Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
+Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 LoRA-PureMoE 50.4 37.6 22.3 13.8 23.1 14.1 31.9 21.8 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
 As shown in Table 8, we compared two settings: PureMoE without specifying modality experts and Random Modality-Specific MoE with randomly assigned experts.
@@
 However, if some crucial text experts were chosen as audio experts, the initial LLM suffers substantial damage, resulting in noticeable performance degradation.
+LoRA-PureMoE builds on PureMoE by applying LoRA-based fine-tuning to the LLM component.
+Although this mitigates the degradation of the LLM’s text capabilities, the speech modality learning remains insufficient.
 Using an adaptive modality assignment method effectively selects which experts are suitable as audio experts and text experts, achieving better performance.
-Appendix D Border Impact
-DeepTalk leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs.
+B.2 Broader Impact
+DeepOmni leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs.
 These limitations include the potential to generate inaccurate information or biased outputs.
 To address these issues, we enhanced the model’s speech interaction capabilities using AudioQA and performed speech-language instruction tuning on high-quality datasets.
-Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepTalk model in any downstream applications.
+Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepOmni model in any downstream applications.
+Appendix C The Use of Large Language Models
+We used LLMs for grammar correction and proofreading.